a 2 tonne car is towed by a 2.5 tonne truck which moves off from rest and reaches a speed of 2m/s before the rope tightens on the stationary car. the rope then stretches taking up the strain for 0.15 seconds. find the common speed when the rope ceases to stretch
1
Expert's answer
2020-02-11T10:11:35-0500
"Mu=(m+M)v"

"(2.5)(2)=(2+2.5)v"

"v=1.1\frac{m}{s}"